OpenStudy (happyfeet8040):

Anyone? Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. y = -x + 6 y = x - 6 y = -x - 2 y = x + 2

rishavraj (rishavraj):

the line is perpendicular to y = -x + 6 so its slope will be 1.... all bcoz if two lines are perpendicular then product of their slope is -1

OpenStudy (anonymous):

More generically, the slope of a line perpendicular to another line will be the negative reciprocal. For example: If the slope of one line is 8, the line perpendicular to it will have a slope of -1/8. Now that we know that the slope of the line perpendicular to y = -x + 6 is 1, we can solve for the y-intercept by plugging in the the point (-4, -2) into y = x +b and solving for b.

OpenStudy (happyfeet8040):

So, C?

OpenStudy (happyfeet8040):

Or A

OpenStudy (happyfeet8040):

@dcabeche

OpenStudy (happyfeet8040):

x+b would be 6 after you plug it in

OpenStudy (happyfeet8040):

Am I right?

OpenStudy (anonymous):

The equation that you are solving is \[-2 = -4 + b\] \[b = 2\] Therefore, the equation of our line is \[y = x + 2\]

OpenStudy (happyfeet8040):

Oh, okay. Gotcha. Thanks. :)